# SupertonicTTS: Towards Highly Scalable and Efficient Text-to-Speech System

### v2

@@ A.2.1 Reference encoder
The reference encoder is composed of a linear layer, 6 ConvNeXt blocks, and 2 cross-attention layers. The linear layer transforms temporally compressed latents with a dimension of 144 to hidden representations with a dimension of 128. The kernel size and intermediate dimension of all ConvNeXt blocks are set to 5 and 512, respectively. In the cross-attention layers, three linear layers with the same input and output dimensions are used to generate query, key, and value. To obtain a fixed number of vectors (i.e., the reference value shown in Figure 4 (a)) representing reference speech, 50 learnable vectors with a dimension of 128 are used in the first attention block.
@@ A.2.2 Text encoder
The text encoder consists of an embedder, 6 ConvNeXt blocks, 4 self-attention blocks, and 2 cross-attention layers. The embedder maps each character to a 128-dimensional vector with a simple lookup table. The kernel size and intermediate dimension of ConvNeXt blocks are set to 5 and 512, respectively. The self-attention blocks follow the transformer encoder architecture, configured with 512 filter channels, 4 attention heads, and rotary position embedding. The cross-attention layers consist of three linear layers with same input and output dimensions, and the first cross-attention layer utilizes 50 learnable vectors (i.e., the reference key shown in Figure 4 (b)), each with a dimension of 128. These 50 vectors are reused as keys in the VF estimator.
@@ A.2.3 VF estimator
The first linear layer in the VF estimator maps 144-dimensional noisy latents to 256-dimensional hidden representations. The main block, highlighted with dotted lines in Figure 4 (c), is composed of 4 dilated ConvNeXt blocks, 2 standard ConvNeXt blocks, TimeCondBlock, TextCondBlock, and RefCondBlock. Each ConvNeXt block has a kernel size of 5 and an intermediate dimension of 1024. The dilation rates for the four dilated ConvNeXt blocks are set to 1, 2, 4, and 8, respectively. TimeCondBlock employs a single linear layer to project a 64-dimensional time embedding onto the channel dimension of input and performs time conditioning via global addition. The time embedding is computed using the same method as in Grad-TTS [47]. TextCondBlock and RefCondBlock employ a cross-attention mechanism to incorporate text and reference speech information, respectively. This structure is repeated four times (). Finally, 4 additional ConvNeXt blocks are applied, followed by a linear layer that maps the 256-dimensional representation back to a 144-dimensional output.
@@ A.3 Duration predictor
We illustrate the detailed architecture of the duration predictor in Figure 5. Since predicting total duration is simpler than phoneme-level duration prediction, we designed this module to be lightweight. Notably, it contains only about 0.5 million parameters.
@@ A.3.2 DP text encoder
The DP text encoder comprises an embedder, 6 ConvNeXt blocks, 2 self-attention blocks, and a linear layer. The embedder converts character-level text input into 64-dimensional vectors. The kernel size and intermediate dimension of each ConvNeXt block are set to 5 and 256, respectively. A learnable 64-dimensional vector, referred to as the utterance token in Figure 5 (b), is prepended to the output of the ConvNeXt blocks. The self-attention blocks have 256 filter channels, 2 attention heads, and incorporate rotary position embeddings. Finally, the first vector from the output of the last self-attention block passes through a linear layer with the same input and output dimension, producing an utterance-level text embedding.
@@ A.3.3 Duration estimator
The duration estimator consists of two linear layers with a PReLU activation. The first layer maintains the input and output dimensions at 164, while the second layer maps the output to a single scalar value. The outputs from the DP reference encoder and the DP text encoder are concatenated before being passing to the first linear layer, as shown in Figure 5 (c).
@@ Appendix D Additional experimental results
D.1 Experimental results on evaluation of context-sharing batch expansion
@@
Figure 6 presents validation loss and ASR results throughout the entire training process. Increasing the expansion factor consistently accelerates convergence for both validation loss and WER. In contrast, while increasing the batch size helps reduce validation loss, it slows the convergence of WER. This indicates that simply increasing the batch size is not sufficient for achieving accurate text-speech alignment. Additionally, this experiment demonstrates that the proposed batch expansion algorithm not only accelerates loss convergence but also alleviates issues such as word skipping, repetition, and mispronunciation.
D.2 Perceptual evaluation of speech quality
To evaluate the perceptual quality of synthesized speech, we conducted subjective listening tests via Amazon Mechanical Turk (MTurk). Since the baseline models (VALL-E, VoiceBox, Mega-TTS 2, CLaM-TTS, and DiTTo-TTS) are not officially open-sourced, we resorted to using the audio samples available on their respective demo pages. However, Mega-TTS 2 was excluded from the evaluation because the full-length prompts used for generating its samples were not provided in the demo page. For each of the remaining baselines, we prepared 15 paired samples, where each pair consisted of one sample from the baseline and one from SupertonicTTS. We then asked 30 participants to evaluate each pair based on two criteria: (1) naturalness - which sample sounds more natural, and (2) speaker similarity - which sample sounds more similar to the reference speaker. For the speaker similarity evaluation, the corresponding reference speech was also provided. All audio files were resampled to before presentation. The participants were instructed to select one of five responses. For quantitative analysis, these responses were mapped to numerical values ranging from -2 to +2 (i.e., -2, -1, 0, 1, 2), where positive scores indicate a preference for SupertonicTTS over the baseline. Figure 7 illustrates the survey interface presented to the evaluation participants.
The results of the subjective evaluation are shown in the left side of Figure 8. For naturalness, SupertonicTTS was generally preferred over VALL-E, CLaM-TTS, and DiTTo-TTS, as indicated by the positive preference scores. However, the subjects perceived VoiceBox samples as more natural than those from SupertonicTTS. In terms of speaker similarity, SupertonicTTS was clearly favored over VoiceBox and slightly outperformed the other baselines. The contrasting performance in the results relative to VoiceBox prompted a further investigation. We analyzed the reference speech samples used in the subjective listening test. Specifically, we used TorchAudio-Squim [31] to estimate the perceptual evaluation of speech quality (PESQ) [48] and scale-invariant signal-to-distortion ratio (SI-SDR) of these samples. The results are shown in the right side of Figure 8. The analysis revealed that VoiceBox’s reference speech had significantly lower PESQ and SI-SDR scores than the comparable-quality references for the other baselines. Based on these findings, we conjecture that SupertonicTTS more faithfully reflects a given reference speech than VoiceBox, and the lower quality of VoiceBox’s reference audio likely caused the SupertonicTTS output to be perceived as less natural, despite its strong performance. Overall, these experimental results support that SupertonicTTS achieves competitive performance on par with recent state-of-the-art zero-shot TTS models.
Appendix E Limitations and future work
@@
Appendix F Broader impacts

### v3

@@ A.2.1 Reference encoder
The reference encoder is composed of a linear layer, 6 ConvNeXt blocks, and 2 cross-attention layers. The linear layer transforms temporally compressed latents with a dimension of 144 to hidden representations with a dimension of 128. The kernel size and intermediate dimension of all ConvNeXt blocks are set to 5 and 512, respectively. In the cross-attention layers, three linear layers with the same input and output dimensions are used to generate query, key, and value. To obtain a fixed number of vectors (i.e., the reference value shown in Fig. 4 (a)) representing reference speech, 50 learnable vectors with a dimension of 128 are used in the first attention block.
@@ A.2.2 Text encoder
The text encoder consists of an embedder, 6 ConvNeXt blocks, 4 self-attention blocks, and 2 cross-attention layers. The embedder maps each character to a 128-dimensional vector with a simple lookup table. The kernel size and intermediate dimension of ConvNeXt blocks are set to 5 and 512, respectively. The self-attention blocks follow the transformer encoder architecture, configured with 512 filter channels, 4 attention heads, and rotary position embedding. The cross-attention layers consist of three linear layers with same input and output dimensions, and the first cross-attention layer utilizes 50 learnable vectors (i.e., the reference key shown in Fig. 4 (b)), each with a dimension of 128. These 50 vectors are reused as keys in the VF estimator.
@@ A.2.3 VF estimator
The first linear layer in the VF estimator maps 144-dimensional noisy latents to 256-dimensional hidden representations. The main block, highlighted with dotted lines in Fig. 4 (c), is composed of 4 dilated ConvNeXt blocks, 2 standard ConvNeXt blocks, TimeCondBlock, TextCondBlock, and RefCondBlock. Each ConvNeXt block has a kernel size of 5 and an intermediate dimension of 1024. The dilation rates for the four dilated ConvNeXt blocks are set to 1, 2, 4, and 8, respectively. TimeCondBlock employs a single linear layer to project a 64-dimensional time embedding onto the channel dimension of input and performs time conditioning via global addition. The time embedding is computed using the same method as in Grad-TTS [55]. TextCondBlock and RefCondBlock employ a cross-attention mechanism to incorporate text and reference speech information, respectively. This structure is repeated four times (). Finally, 4 additional ConvNeXt blocks are applied, followed by a linear layer that maps the 256-dimensional representation back to a 144-dimensional output.
@@ A.3 Duration predictor
We illustrate the detailed architecture of the duration predictor in Fig. 5. Since predicting total duration is simpler than phoneme-level duration prediction, we designed this module to be lightweight. Notably, it contains only about 0.5 million parameters.
@@ A.3.2 DP text encoder
The DP text encoder comprises an embedder, 6 ConvNeXt blocks, 2 self-attention blocks, and a linear layer. The embedder converts character-level text input into 64-dimensional vectors. The kernel size and intermediate dimension of each ConvNeXt block are set to 5 and 256, respectively. A learnable 64-dimensional vector, referred to as the utterance token in Fig. 5 (b), is prepended to the output of the ConvNeXt blocks. The self-attention blocks have 256 filter channels, 2 attention heads, and incorporate rotary position embeddings. Finally, the first vector from the output of the last self-attention block passes through a linear layer with the same input and output dimension, producing an utterance-level text embedding.
@@ A.3.3 Duration estimator
The duration estimator consists of two linear layers with a PReLU activation. The first layer maintains the input and output dimensions at 164, while the second layer maps the output to a single scalar value. The outputs from the DP reference encoder and the DP text encoder are concatenated before being passing to the first linear layer, as shown in Fig. 5 (c).
@@ Appendix D Additional experimental results
D.1 Trade-off between the number of function evaluations and synthesis quality
| NFE | RTF ↓ | WER ↓ | SIM ↑ | NISQA ↑ | |
| GT | - | - | 2.181 | 0.677 ± 0.006 | 4.070 ± 0.029 |
| Ours | 4 | 0.006 | 11.43 | 0.335 ± 0.003 | 2.623 ± 0.020 |
[…]
| 128 | 0.140 | 2.693 | 0.468 ± 0.003 | 4.070 ± 0.014 |
SupertonicTTS synthesizes speech through Euler’s method during inference. By adjusting the number of function evaluations (NFE), we can balance the trade-off between synthesis speed and quality. To quantify this relationship, we generated five samples per utterance from the LS-clean by varying NFE values while keeping a CFG coefficient to 3. These samples were then evaluated in terms of RTF, WER, NISQA score [45], and speaker similarity (SIM). SIM was calculated as the cosine similarity between speaker embeddings from the generated speech and the corresponding 3-second reference, using the WavLM-TDNN model [6].
The evaluation results, presented in Table 8, demonstrate a general trend that increasing NFE leads to improvements in the WER, SIM, and NISQA scores. Specifically, the best scores for WER and SIM were obtained at NFE values of 32 and 16, respectively. This suggests that NFE values exceeding 32 effectively capture intelligibility, prosodic naturalness, and speaker identity. Meanwhile, NISQA scores exhibited a consistent positive correlation with NFE, suggesting that higher NFE values result in enhanced audio fidelity. However, this improvement comes at the cost of increased processing time, as reflected in the RTF. Given this trade-off, we select as the optimal balance between quality and efficiency.
D.2 Experimental results on evaluation of context-sharing batch expansion
@@
Fig. 6 presents validation loss and ASR results throughout the entire training process. Increasing the expansion factor consistently accelerates convergence for both validation loss and WER. In contrast, while increasing the batch size helps reduce validation loss, it slows the convergence of WER. This indicates that simply increasing the batch size is not sufficient for achieving accurate text-speech alignment. Additionally, this experiment demonstrates that the proposed batch expansion algorithm not only accelerates loss convergence but also alleviates issues such as word skipping, repetition, and mispronunciation.
Appendix E Subjective listening test
Fig. 7 illustrates the survey interface presented to the evaluation participants. All audio files were resampled to before the subjective listening tests.
Appendix F Public datasets
Table 9 provides a list of public datasets used for training the speech autoencoder.
Appendix G Limitations and future work
@@
Appendix H Broader impacts